OpenStudy (anonymous):

find values of x for which f(x)=g(x) f(x)=x^2+2x+1, g(x)=3x+3

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

can you show me the steps please?

OpenStudy (anonymous):

all you need to do is solve x^2+2x+1 = 3x+3 put it in standard form: x^2-x-2 = 0 subtracted 3x+3 from both sides and simplified. (x+1)(x-2)=0 factored you can take it from here. the answer is now obvious.

sam (.sam.):

Equate them together and solve since f(x)=g(x) x^2+2x+1=3x+3

OpenStudy (anonymous):

so the values would be like 1 and -2?? im confused.

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

no, from here, (x+1)(x-2)=0 you set each brackets =0, (x+1)=0, (x-2)=0 then solve for 2 values of x

OpenStudy (anonymous):

ohh so -1 and 2 thank you! :)

sam (.sam.):

yes

OpenStudy (anonymous):

thanks sam! :)

sam (.sam.):

welcome :)